for reaction a+2b gives c+d if the initial moles of a and b are 2 and 4 in 1 litre then find out Kc if total moles at equilibrium are 5

Grade:11

1 Answers

Khushi
6 years ago
a+2b gives c+dMolesey taken initially 2 4 0 0 Moles at equilibrium (2-x) (4-2x) x x Total moles (2-x) +(4-2x) +2x = 5x=1Kc= 1×1upon 1×2 = 0.5 Kc= 0.5